Define h(0,0) in such a way that h is continuous at origin h(u,v)=ln⁡((nu^2-u^2 v^2+nv^2)/(u^2+v^2 ))



1
Expert's answer
2021-05-11T12:39:36-0400

"Given, h(u,v)=ln\u2061\\frac{nu^2\u2212u^2v^2+nv^2}{u^2+v^2\u200b}.\\newline\n\\text{Solve for h,}\\newline\n\nh(u,v)=ln\u2061\\dfrac{n(u^2+v^2)-(uv)^2}{u^2+v^2 }\\newline\n=\\ln\u2061(\\dfrac{n(u^2+v^2)}{u^2+v^2 }-\\frac{(uv)^2}{u^2+v^2 })\\newline\n=\\ln\u2061(n-\\frac{(uv)^2}{u^2+v^2 })\\newline\n\\text{ Taking limit.}(u,v)\\to(0,0)\\newline\nlim_{(u,v)\\to(0,0)} h(u,v)\\newline\n=lim_{(u,v)\\to(0,0)} \\ln\u2061(n-\\frac{(uv)^2}{u^2+v^2 })\\hspace{1cm}(\\frac{0}{0} form)\\newline\n(put \\space v=mu.)\\newline\n=lim_{u\\to0} \\ln\u2061(n-\\frac{u^4m^2}{u^2(1+m^2) })\\newline\n=lim_{u\\to0} \\ln\u2061(n-\\frac{u^2}{(1+m^2) })\\newline\n\n=ln(n)\\newline\n\\text{Thus,} h(x,y)=ln(n), at (x,y)=(0,0)."